OpenStudy (brynndelyn):

Which of the following is the simplified form of [5sqrt{x}]^4 x^1/5 x^4/5 x^4/20 x

rishavraj (rishavraj):

is it \[(\sqrt[5]{x})^4\]

rishavraj (rishavraj):

if it is so then u know \[(\sqrt[a]{x})^b = x ^{\frac{ b }{ }}\]

OpenStudy (brynndelyn):

Thanks!

rishavraj (rishavraj):

u r most welcome...:)